Firms HL and LL are identical except for their financial leverage ratios and the interest rates they pay on debt. Each has $20 million in invested capital, has $3 million of EBIT, and is in the 40% federal-plus-state tax bracket. Firm HL, however, has a debt-to-capital ratio of 55% and pays 13% interest on its debt, whereas LL has a 20% debt-to-capital ratio and pays only 8% interest on its debt. Neither firm uses preferred stock in its capital structure.

  1. Calculate the return on invested capital (ROIC) for each firm. Round your answers to two decimal places.

    ROIC for firm LL is ___%
    ROIC for firm HL is ___%

  2. Calculate the rate of return on equity (ROE) for each firm. Round your answers to two decimal places.

    ROE for firm LL is ___%
    ROE for firm HL is ___%

  3. Observing that HL has a higher ROE, LL's treasurer is thinking of raising the debt-to-capital ratio from 20% to 60% even though that would increase LL's interest rate on all debt to 15%. Calculate the new ROE for LL. Round your answer to two decimal places.

    ___%
Answer #1

(a) ROIC= Net operating profit after tax/Invested capital

NOPAT= EBIT(1-Tax rate) = 3 million*0.60= 1.80 million

ROIC= (1.80/20) = 9%

It is same for both the companies as EBIT and invested capital of both the companies is same and tax rate is also the same.

(b) ROE= Return available for equity shareholders/Equity capital

Firm LL has 80% equity capital i.e. 16 million

Return for equity= (EBIT-interst)*(1-tax rate)

=(3-0.32)*0.6 =1.608

ROE= 1.608/16= 10.05%

Firm HL has 45% equity capital= 9 million

Return for equity= (EBIT-interst)*(1-tax rate)

=(3-0.1.43)*0.60= 0.942

ROE= 0.942/9= 10.47%

(c) New ROE of LL

Return for equity =(3-1.80)*0.6= 0.72

Equity capital- 40% of 20 million= 8 million

New ROE= 0.72/8= 9%